# Optimizing the pump coupling for a three-wave mixing Josephson parametric amplifier

### v1

@@ III Realization of a filter-coupled SNAIL parametric amplifier
Figure 3: Filter-coupled JPA in which the Josephson dipole is a SNAIL array. (a) Schematic of a SNAIL parametric amplifier with two filters as coupling networks, with the signal flow graphs representing the desired scattering properties at signal (red) and pump (blue) frequencies respectively. The frequency landscape is sketched in the top right. (b) Optical image of a microstrip device designed to achieve the indicated scattering properties, with an SEM image of the array of 10 SNAILs.
@@
In the last section, the goals and requirements for the circuit synthesis are formulated in terms of the scattering parameters of a 3-port linear network. However, it is technically challenging to directly synthesize this 3-port network in distributed-element circuit. Therefore, as illustrated in Fig. 3(a), we separate the design into two 2-port networks – a pump-port network and a signal-port network, with an array of SNAILs[28] embedded in between, playing the role of the Josephson dipole. The pump-port network prevents transmission at signal frequency $\omega_{s}$ while providing near unity transmission at pump frequency $\omega_{p}$. Conversely, the signal-port network prevents transmission at $\omega_{p}$ while allowing partial transmission at $\omega_{s}$ with a rate $\kappa_{0}$, which is critical in forming the resonant mode of the amplifier. The signal-port network also suppresses pump leakage in addition to the suppression originating from the improved pump coupling efficiency according to Eq. (3).
@@
Instead of the standard two-port microwave filter synthesis process based on scattering parameters[29, 30], we formulated the design goals using the output impedances $Z^{(s)}$ and $Z^{(p)}$, viewed from each end of the Josephson dipole towards the signal and pump ports as shown in Fig. 3(a), respectively. This allows the resonance condition for the concatenated circuit to be represented as an additional constraint on the sum of the two impedances $Z^{(s)}+Z^{(p)}=Z^{\mathrm{Th}}$ according to Eq. (5). We discuss in Appendix B how to bridge between the output impedances and the scattering parameters, and how to formulate the resonance condition.
@@
For compatibility with our in-house fabrication and packaging process (Appendix A), we designed and implemented the amplifier in a microstrip architecture. We employed a compact 3-stage hairpin bandpass filter (with the poles near $\omega_{p}$) for the pump-port network and a 3-stage radial-stub lowpass filter for the signal-port network to achieve the desired frequency-selective properties. Specifically, the slow frequency roll-off property of the radial-stub lowpass filter allows us to position the amplifier mode within its transition band, facilitating control of $\kappa_{0}$.
@@
The synthesis process is conducted using AWR Microwave Office. We began by parameterizing the dimensions of the hairpin and radial-stub filter structures in the distributed element circuit. Initial design parameters were chosen to achieve desired filtering performance within the intended signal frequency band (4.8-5.2 GHz) and pump frequency band (9.6-10.4 GHz). In addition, we aimed for a resonant mode with the following properties: (1) resonance frequency $\omega_{0}/2\pi\approx 5$ GHz, (2) linewidth $\kappa_{0}/2\pi\approx 100$ MHz (dominated by external coupling via the signal port), (3) external coupling rate via the pump port below 0.5 MHz. These requirements were formulated into constraints on the output impedances of the filter networks. We tuned the design parameters with the optimizer tool in AWR Microwave Office to simultaneously meet all these objectives. Finally, we laid out the circuit and performed electromagnetic simulations in AWR Axiem to verify and further refine the circuit design.
It is crucial to note that the output impedances $Z^{(s)}[\omega]$ and $Z^{(p)}[\omega]$ are evaluated assuming $50\Omega$ matching at ports 1 and 2. Experimentally ensuring this matching condition at all frequency bands of interest is essential for the design to function in the physical implementation.
@@
IV F-SPA characterization and comparison with a regular SPA
In this section, we report the characterization of the performance of the F-SPA described in the last section. To demonstrate the advantage of the F-SPA, we compare its power efficiency, pump leakage, and noise performance with that of a standard capacitor-coupled SPA (referred to in the following as C-SPA)[8, 15].
@@ IV F-SPA characterization and comparison with a regular SPA
We first measure the linear reflection from port 1 to extract the resonance frequency and linewidth of the F-SPA mode at different external flux bias $\Phi_{\mathrm{ext}}$, as shown in Fig. 4. The measured resonance frequency agrees well with that from the AWR Axiem simulation over the full flux tunable range of the device. The measured linewidth, however, is noticeably lower than that from simulation near zero external flux. This deviation is likely caused by fabrication uncertainty that shifts the pass band of the signal port filter, and impedance mismatch on this port. Due to the low-pass nature of this radial-stub filter, the linewidth of the F-SPA mode is more sensitive to the frequency roll-off of the filter when in the high frequency part of its tunable range.
@@
At each flux bias point, we then apply from port 2 a pump tone at frequency $\omega_{p}=2\omega_{0}$ with the power that generates a small-signal 20 dB phase-preserving gain in reflection. Under this condition, we measure the pump leakage power $P_{\mathrm{leak}}$ (at port 1), output signal 1 dB compression power $P_{1\mathrm{dB}}^{\mathrm{out}}$ and the added noise at the frequency 100 kHz below $\omega_{0}$.
@@
Figure 5: Characterization of the F-SPA (green dots) and C-SPA (orange triangles) with 20 dB reflection gain over their operating frequency ranges. (a) Output signal power at 1 dB compression points versus pump power at device pump port. Positions of the sloped lines indicates power efficiency $\eta_{p}$. (b) Pump leakage power measured at device signal port. (c) Added noise temperature $T_{\mathrm{add}}$, with a black dashed line indicating the standard quantum limit $\hbar\omega/2k_{B}$.
@@
The same characterization of a C-SPA is performed with identical measurement setup. The C-SPA is implemented by embedding the SNAIL array at the current antinode of a half-wave transmission line resonator, with a large coupling capacitance at port 1 and a small coupling capacitance at port 2. For a fair comparison, the resonance frequency, linewidth and 3-wave mixing strength of these two devices are designed to be nominally identical. The resonance frequencies of both the F-SPA and the C-SPA are designed to be between 4.5 GHz and 5.5 GHz such that their pump frequencies, between 9.0 GHz and 11.0 GHz, are within the pass band of the microwave components on the output line. Due to fabrication uncertainty, the linewidth of the C-SPA turned out in the range of 60 to 90 MHz, approximately a factor of two smaller than that of the F-SPA. This would result in about 6 dB lower pump power required for the same 20 dB gain.
@@
As shown in Fig. 5(a), the F-SPA requires 20 to 25 dB less pump power than the C-SPA to generate 20 dB gain across the full operating range. The F-SPA shows higher $P_{1\mathrm{dB}}^{\mathrm{out}}$ due to its larger linewidth compared to the C-SPA. Overall, the F-SPA achieves an average power efficiency of $\eta_{p}=3\times 10^{-3}$, which is 30 dB better than the C-SPA and consistent with the 27 dB improvement predicted by our linear circuit simulation.
@@
The pump leakage power, as plotted in Fig. 5(b), is on average -113 dBm for the F-SPA and -77 dBm for the C-SPA. This dramatic 36 dB improved suppression in pump leakage on the F-SPA originates from the combination of the improved pump coupling efficiency and the reflectivity from the signal-port filter.
@@
It is worth pointing out that the pump for a typical 3-wave-mixing JPA could fall outside the band of the circulators between the signal source and the JPA. The out-of-band reverse isolation of the circulators are generally not optimized. It is therefore of practical advantage to suppress pump leakage at the device level instead of relying on external components.
@@
The added noise temperature (with respect to the input signal) of these two amplifiers are shown in Fig. 5(c). The added noise and output power of the amplifiers are calibrated using a shot noise tunnel junction (SNTJ)[31] as explained in Appendix A. Both devices have added noise that are at most twice the quantum limit for phase-preserving amplification[32]. The excess added noise is consistent for both devices, which is most likely due to residual thermal noise from the signal input line. Thermal noise from the pump line, on the other hand, does not contribute to added noise in these measurements as we will discuss in the next section.
@@ V Robustness against pump port thermal noise on the F-SPA
The F-SPA requires significantly less pump power to operate compared to the C-SPA, leading to a reduced heat load to the dilution refrigerator from power dissipation on the pump line cryogenic attenuators. In this section, we investigate the effect of thermal noise from the pump line on these two devices. Remarkably, we demonstrate that the F-SPA maintains its near quantum-limited noise performance even when exposed to 4 K noise from the pump port. This characteristic affords an opportunity for further heat load reduction by decreasing the pump line attenuation at the millikelvin stage of the refrigerator.
@@
To achieve quantum-limited noise performance on an amplifier, the signal port input noise temperature $T^{\mathrm{in},1}$ must be thermalized to the quantum noise level. The quantum limit of half quanta added noise for a phase-preserving parametric amplifier originates from the input quantum noise at the idler frequency ($\omega_{i}=\omega_{p}-\omega_{s}$). Having another port for pump power delivery, however, introduces additional channels for added noise.
@@
To quantify the effect of pump port input noise on amplifier noise performance, we inject noise along with the coherent pump tone into the pump line and measure the output noise of the amplifier, as illustrated in Fig. 6(a). The injected noise is produced by mixing an artificially generated noise, which has a flat-top power spectrum density from DC to 1 GHz, with a local oscillator near signal and pump frequency, respectively. In this way, we separately investigate the effects of the noise in the signal and idler frequency bands (referred to as signal-band noise, the idler frequency for a degenerate amplifier is usually close to the signal frequency) and the noise in the pump frequency band (referred to as pump-band noise) respectively. The output noise power spectral density is measured at the frequency $\omega_{s}$ of 20 dB small-signal phase-preserving gain. Both input and output noise power spectral densities are calibrated into noise temperatures $T_{N}[\omega]$ at the device ports (Appendix A). We perform this experiment on the F-SPA at $\omega_{s}/2\pi=5.04$ GHz and the C-SPA at $\omega_{s}/2\pi=4.76$ GHz, which corresponds to the operating points with highest $P_{\mathrm{1dB}}$ for the two devices respectively.
@@
Without noise injection, the output noise of both the F-SPA (marked circle) and the C-SPA (marked triangle) are close to the quantum limit, given by $T^{\mathrm{out},1}_{N}[\omega_{s}]=G\cdot\hbar\omega_{s}/k_{B}$. Upon injecting noise in the signal and idler frequency bands, the output noise temperature increases linearly with the pump-port input noise temperature. As indicated by the yellow curved lines in Fig. 6(b), the F-SPA is characterized by a fitted linear coefficient of 0.30, while the C-SPA exhibits a higher coefficient of 3.6. This coefficient should equal to twice the signal transmission coefficient defined in Eq.8 due to reciprocity of transmission at both the signal and idle frequencies. Compared to the C-SPA, the F-SPA exhibits an order of magnitude better suppression against signal-band noise from the pump line. Notably, the F-SPA maintains near quantum-limited performance even with noise injection up to 4 K, showing only a minimal increase in added noise amounting to 0.05 quanta. This robust performance indicates that the thermalization requirement for the pump line of F-SPA can be significantly relaxed compared to conventional designs.
@@
Figure 6: Pump line noise injection experiment. (a) Experimental setup: output noise temperature from the signal port is measured as a function of the pump port input noise temperature, under the 20 dB gain pumping condition. (b) Data obtained from F-SPA (circles) and C-SPA (triangles) under injected noise in signal and idler frequency bands (yellow) and pump frequency band (blue) respectively. Yellow lines represent linear fits to the data, with fitted slope 3.6 for the C-SPA and 0.30 for the F-SPA. The black horizontal line indicates the standard quantum limit (SQL) of the added noises.
@@
When injecting noise at pump frequency, the output noise temperature of the F-SPA starts to increase drastically (i.e. beyond linear dependence) at a pump-band noise temperature of $10^{3}$ K, as shown in  6(b). In contrast, the output noise temperature of the C-SPA barely starts to increase at $10^{5}$ K of pump-band noise. This 20 dB difference in the threshold pump-band noise temperature is just another manifestation of the improved pump coupling efficiency on the F-SPA. However, the exact mechanism behind how pump-band noise cause signal-bands output noise to increase requires further investigation.
@@
Improving the pump coupling efficiency indeed increases the susceptibility of F-SPA to noise at the pump frequency. However, this should not degrade its noise performance in standard applications where the pump-band noise temperature typically remains well below 1000 K. On the other hand, the filtering at the pump port of the F-SPA provides improved suppression of signal-band noise, thereby offering an opportunity for substantial heat load reduction. This reduction may be achieved through decreasing or potentially eliminating attenuation at the millikelvin stage of the dilution refrigerator.

### v2

@@ III Realization of a filter-coupled SNAIL parametric amplifier
Figure 3: Filter-coupled JPA in which the Josephson dipole is a SNAIL array. (a) Schematic of a SNAIL parametric amplifier with two filters as coupling networks, with the signal-flow graphs representing the desired scattering properties at the signal (red) and pump (blue) frequencies respectively. The frequency landscape is sketched in the top right. (b) Optical image of a microstrip device designed to achieve the indicated scattering properties, with a scanning electron microscope image of the array of 10 SNAILs.
@@
In the last section, the goals and requirements for the circuit synthesis are formulated in terms of the scattering parameters of a 3-port linear network. However, it is technically challenging to directly synthesize this 3-port network in distributed-element circuit. Therefore, as illustrated in Fig. 3(a), we separate the design into a pair of 2-port networks — a pump-port network and a signal-port network — with an array of SNAILs Frattini et al. (2017) embedded in between, playing the role of the Josephson dipole. The pump-port network prevents transmission at the signal frequency $\omega_{s}$ while providing near-unity transmission at pump frequency $\omega_{p}$. Conversely, the signal-port network prevents transmission at $\omega_{p}$ while allowing partial transmission at $\omega_{s}$ with a rate $\kappa_{0}$, which is critical in forming the resonant mode of the amplifier. The signal-port network also suppresses pump leakage, in addition to the suppression originating from the improved pump coupling efficiency according to Eq. (4).
@@
Instead of the standard two-port microwave filter synthesis process based on scattering parameters Pozar (2005); Matthaei et al. (1980), we formulated the design goals using the output impedances $Z^{(s)}$ and $Z^{(p)}$, viewed from each end of the Josephson dipole toward the signal and pump ports, respectively, as shown in Fig. 3(a). This allows the resonance condition for the concatenated circuit to be represented as an additional constraint on the sum of the two impedances $Z^{(s)}+Z^{(p)}=Z^{\mathrm{Th}}$ according to Eq. (5). We discuss in Appendix B how to bridge between the output impedances and the scattering parameters, and how to formulate the resonance condition.
@@
For compatibility with our in-house fabrication and packaging process (Appendix A), we designed and implemented the amplifier in a microstrip architecture. We employed a compact three-stage hairpin bandpass filter (with the poles near $\omega_{p}$) for the pump-port network and a three-stage radial-stub lowpass filter for the signal-port network to achieve the desired frequency-selective properties. Specifically, the slow frequency roll-off property of the radial-stub lowpass filter allows us to position the amplifier mode within its transition band, facilitating control of $\kappa_{0}$.
@@
The synthesis process was conducted using AWR Microwave Office. We parametrized the hairpin and radial-stub structures based on our target signal (4.8-5.2 GHz) and pump (9.6-10.4 GHz) frequency bands. In addition, we aimed for a resonant mode at $\omega_{0}/2\pi\approx 5$ GHz with linewidth $\kappa_{0}/2\pi\approx 100$ MHz dominated by signal-port damping. Formulating these requirements as constraints on the output impedances of the filter networks, we optimized the design parameters and verified the layout through electromagnetic simulations in AWR Axiem.
The output impedances $Z^{(s)}[\omega]$ and $Z^{(p)}[\omega]$ are evaluated assuming $50\Omega$ matching at ports 1 and 2. Ensuring this matching condition experimentally across all relevant frequency bands is important for proper implementation, which requires attention, as pump frequencies often exceed the operational band of external microwave components.
@@
IV F-SPA characterization and comparison with a conventional SPA
In this section, we report the characterization of the performance of the F-SPA described in the last section, and demonstrate its advantages over an SPA of conventional design Frattini et al. (2018); Sivak et al. (2019).
@@ IV F-SPA characterization and comparison with a conventional SPA
We first measure the linear reflection from port 1 to extract the resonance frequency and linewidth of the F-SPA mode at different external flux bias values $\Phi_{\mathrm{ext}}$, as shown in Fig. 4. The measured resonance frequency agrees well with that from the AWR Axiem simulation over the full flux tunable range of the device. The measured linewidth, however, is noticeably lower than that from the simulation near zero external flux. This deviation is likely caused by fabrication uncertainty, which shifts the pass band of the signal-port filter, and impedance mismatch on this port. Due to the lowpass nature of this radial-stub filter, the linewidth of the F-SPA mode is more sensitive to the frequency roll-off of the filter when in the high frequency part of its tunable range.
@@
At each flux bias point, from port 2, we then apply a pump tone at frequency $\omega_{p}=2\omega_{0}$ with the power that generates a small-signal 20-dB phase-preserving gain in reflection. Under this condition, we measure the pump leakage power $P_{\mathrm{leak}}$ (at port 1), output signal 1-dB compression power $P_{1\mathrm{dB}}^{\mathrm{out}}$ and the added noise at the frequency 100 kHz below $\omega_{0}$.
@@
Figure 5: Characterization of the F-SPA (green dots) and C-SPA (orange triangles) with 20 dB reflection gain over their operating frequency ranges. (a) Output signal power at 1-dB compression points versus pump power at device pump port. Positions of the sloped lines indicates power efficiency $\eta_{p}$. (b) Pump leakage power measured at device signal port. (c) Added noise temperature $T_{\mathrm{add}}$, with a black dashed line indicating the standard quantum limit (SQL) $\hbar\omega/2k_{B}$.
To demonstrate the advantage of the F-SPA, we compare its power efficiency, pump leakage, and noise performance with that of a control device using capacitive coupling at both ports Sivak et al. (2019). The latter consists of an array of SNAILs embedded at the center of a half-wave transmission line resonator, with a large coupling capacitance at the signal port and a small coupling capacitance at the pump port (see the schematics in the legend of Fig. 5). We referred to this device as the capacitor-coupled SPA or C-SPA.
@@
For a fair comparison, the resonance frequency, linewidth and three-wave mixing strength of the C-SPA and the F-SPA are designed to be nominally identical. The resonance frequencies of both devices are designed to be between 4.5 GHz and 5.5 GHz such that their pump frequencies, between 9.0 GHz and 11.0 GHz, are within the pass band of the microwave components on the output line. Due to fabrication uncertainty, the linewidth of the C-SPA turned out in the range of 60 to 90 MHz, approximately a factor of two smaller than that of the F-SPA. This would result in about 6 dB lower pump power required for the same 20-dB gain.
@@
We perform identical characterization measurements on the two amplifiers using the same experimental setup. As shown in Fig. 5(a), the F-SPA requires 20 to 25 dB less pump power than the C-SPA to generate 20 dB gain across the full operating range. The F-SPA shows higher $P_{1\mathrm{dB}}^{\mathrm{out}}$ due to its larger linewidth compared to the C-SPA. Overall, the F-SPA achieves an average power efficiency of $\eta_{p}=3\times 10^{-3}$, which is 30 dB better than the C-SPA and consistent with the 27-dB improvement predicted by our linear circuit simulation.
@@
The pump leakage power, as plotted in Fig. 5(b), is on average -113 dBm for the F-SPA and -77 dBm for the C-SPA. This dramatic 36-dB improvement in suppression of pump leakage on the F-SPA originates from the combination of the improved pump coupling efficiency and the reflectivity from the signal-port filter.
@@
It is worth pointing out that the pump for a typical three-wave-mixing JPA could fall outside the band of the circulators between the signal source and the JPA. The out-of-band reverse isolation of the circulators is generally not optimized. It is therefore of practical advantage to suppress pump leakage at the device level instead of relying on external components.
@@
The added noise temperatures (with respect to the input signal) of these two amplifiers are shown in Fig. 5(c). The added noise and output power of the amplifiers are calibrated using a shot-noise tunnel junction (SNTJ) Malnou et al. (2024) as explained in Appendix A. Both devices have added noise that is at most twice the quantum limit for phase-preserving amplification Caves (1982). The excess added noise is consistent for both devices, which is most likely due to residual thermal noise from the signal input line. Thermal noise from the pump line, however, does not contribute to added noise in these measurements as we will discuss in the next section.
@@ V Robustness against pump port thermal noise on the F-SPA
The F-SPA requires significantly less pump power to operate compared to the C-SPA, leading to a reduced heat load to the dilution refrigerator from power dissipation on the pump line cryogenic attenuators. In this section, we investigate the effect of thermal noise from the pump line on these two devices. Remarkably, we demonstrate that the F-SPA maintains its near quantum-limited noise performance even when exposed to 4 K noise from the pump port. This characteristic affords an opportunity for further heat load reduction by decreasing the pump-line attenuation at the millikelvin stage of the refrigerator.
@@
To achieve quantum-limited noise performance on an amplifier, the signal port input noise temperature $T^{\mathrm{in},1}$ must be thermalized to the quantum noise level. The quantum limit of half a quantum of added noise for a phase-preserving parametric amplifier originates from the input quantum noise at the idler frequency ($\omega_{i}=\omega_{p}-\omega_{s}$). Having another port for pump power delivery, however, introduces additional channels for added noise.
@@
To quantify the effect of pump port input noise on amplifier noise performance, we inject noise along with the coherent pump tone into the pump line and measure the output noise of the amplifier, as illustrated in Fig. 6(a). The injected noise is produced by mixing artificially generated noise — which has a flat-top power spectral density from DC to 1 GHz — with a local oscillator near signal and pump frequency, respectively. In this way, we separately investigate the effects of the noise in the signal and idler frequency bands (referred to as signal-band noise, the idler frequency for a degenerate amplifier is usually close to the signal frequency) and the noise in the pump frequency band (referred to as pump-band noise). The output noise power spectral density is measured at the frequency $\omega_{s}$ of 20 dB small-signal phase-preserving gain. Both the input and output noise power spectral densities are calibrated into noise temperatures $T_{N}[\omega]$ at the device ports (Appendix A). We perform this experiment on the F-SPA at $\omega_{s}/2\pi=5.04$ GHz and on the C-SPA at $\omega_{s}/2\pi=4.76$ GHz, corresponding to the operating points with highest $P_{\mathrm{1dB}}$ value for each device.
@@
Without noise injection, the output noise of both the F-SPA (marked with circles) and the C-SPA (marked with triangles) are close to the quantum limit, given by $T^{\mathrm{out},1}_{N}[\omega_{s}]=G\cdot\hbar\omega_{s}/k_{B}$. Upon injecting noise into the signal and idler frequency bands, the output noise temperature increases linearly with the pump-port input noise temperature. As indicated by the yellow curved lines in Fig. 6(b), the F-SPA is characterized by a fitted linear coefficient of 0.30, while the C-SPA exhibits a higher coefficient of 3.6. This coefficient should equal to twice the signal transmission coefficient defined in Eq.12 due to reciprocity of transmission at both the signal and idle frequencies. Compared to the C-SPA, the F-SPA exhibits an order of magnitude better suppression against signal-band noise from the pump line. Notably, the F-SPA maintains near quantum-limited performance even with noise injection up to 4 K, showing only a minimal increase in added noise amounting to 0.05 quanta. This robust performance indicates that the thermalization requirement for the pump line of F-SPA can be significantly relaxed compared to conventional designs.
@@
Figure 6: Pump-line noise injection experiment. (a) Experimental setup: output noise temperature from the signal port is measured as a function of the pump-port input noise temperature, under the 20 dB gain pumping condition. (b) Data obtained from F-SPA (circles) and C-SPA (triangles) under injected noise in signal and idler frequency bands (yellow) and the pump frequency band (blue) respectively. Yellow lines represent linear fits to the data, with a fitted slope of 3.6 for the C-SPA and 0.30 for the F-SPA. The black horizontal line indicates the standard quantum limit (SQL) of the added noises.
@@
When injecting noise at pump frequency, the output noise temperature of the F-SPA starts to increase drastically (i.e. beyond linear dependence) at a pump-band noise temperature of $10^{3}$ K, as shown in  6(b). In contrast, the output noise temperature of the C-SPA barely starts to increase at $10^{5}$ K of pump-band noise. This 20 dB difference in the threshold pump-band noise temperature is just another manifestation of the improved pump coupling efficiency on the F-SPA. However, the exact mechanism behind how pump-band noise causes signal-band output noise to increase requires further investigation.
@@
Improving the pump coupling efficiency indeed increases the susceptibility of F-SPA to noise at the pump frequency. However, this should not degrade its noise performance in standard applications where the pump-band noise temperature typically remains well below 1000 K. Nonetheless, the filtering at the pump port of the F-SPA provides improved suppression of signal-band noise, thereby offering an opportunity for substantial heat load reduction. This reduction may be achieved through decreasing or potentially eliminating attenuation at the millikelvin stage of the dilution refrigerator.